OpenStudy (anonymous):

i am confused and need help will medal Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (anonymous):

Since you are given answers, you could check them until you find the right one. For example, to check the first one, do 1 - 2*12 = 1 - 24 = -23. That's not equal to 12, so the first answer cannot be right.

OpenStudy (anonymous):

ooooooooh ok i think i get it know thanks you are a big help

OpenStudy (anonymous):

Joel is right, but if you want to solve it rather than just checking answers, eliminate the X from the equations by adding them together, leading to y=23, and then solve for x from that and one of the equations